OpenStudy (anonymous):

The pressure of earth's atmosphere at sea level is approximately 1000 millibars and decreases exponentially with elevation. At an elevation of 30,000ft, the pressure is one-third of the sea level pressure. At what elevation is the pressure half of the sea-level pressure? At what elevation is it 1% of the sea-level pressure?

OpenStudy (anonymous):

@electrokid @abb0t

OpenStudy (anonymous):

decreases exponentially. so, let the pressure at a height "h" be expressed as: \[P(h)=P_0e^{-kh}\] where P0 = pressure at height h=0 and k = an exponential const. first obtain k using P(30,000)=1000/3

OpenStudy (anonymous):

then, find "h" such that P(h)=1000*(1/100)

OpenStudy (anonymous):

we can re-write our function as: \[ -kh = \ln(P/P_0)\implies kh=\ln(P_0/P)\\ k(30000)=\ln(3)\implies\text{find k}\\ (?)h=\ln(100) \]